OpenStudy (anonymous):

Simplify and write answer in exponential form: 3^3 + 2 • 3^3

OpenStudy (anonymous):

\[3\times3\times3+2\times3\times3\times3\]

OpenStudy (anonymous):

My professor's answer was 3 x 3^3 but idk how she got that

OpenStudy (tkhunny):

Factor this: A + 2A

OpenStudy (anonymous):

3A

OpenStudy (tkhunny):

Now try \((3^{3} + 2\cdot 3^{3})\) Hint: It is EXACTLY the same.

OpenStudy (anonymous):

I'm sorry, but I still don't understand this.. :(

OpenStudy (anonymous):

one apple and two apples are three apples one \(3^3\) and two \(3^3\) are three \(3^3\) and \(3\times 3^3=3^{1+3}=3^4\)

OpenStudy (anonymous):

if your professor really wrote \[3\times 3^3\] he/she was not paying attention

OpenStudy (zehanz):

Try to see it this way: you have: 3^3 + 2 • 3^3, so this is 1 • 3^3 + 2 • 3^3 = 3 • 3^3 (you have 3 times 3^3). Last step: 3 • 3^3 = 3 • 3•3•3 = 3^4 (=81, I fail to see why you would stop after writing it as 3^4, just calculate the outcome...)

OpenStudy (anonymous):

Omg! Thank you! That helped :)

OpenStudy (zehanz):

YW!